OpenStudy (anonymous):

A car travels 100 meters in 5 seconds. What is the speed of the car? speed = distance over time; 1 mile = 1609 meters

OpenStudy (anonymous):

i need a walk through

OpenStudy (michele_laino):

by definition, the speed of the car is: \[\Large speed = \frac{{{\text{distance}}}}{{{\text{time}}}}\]

OpenStudy (anonymous):

100 miles is distance and

OpenStudy (anonymous):

5 secs is time

OpenStudy (anonymous):

So how much is 100m in miles?? 100/1609 right? :D And how much is 5 seconds in hours? 5/3600 right? :D So miles/hour = \[\frac{ 100 }{ 1609 }*\frac{ 3600 }{ 5 }\] 44.7 miles

OpenStudy (campbell_st):

if you multiply both values by 12 you will know how many metres in 1 minute this is because 12 x 5 = 60 then multiply the answer by 60 to find how many metres in 1 hour does that make sense..?

OpenStudy (michele_laino):

\[\Large speed = \frac{{{\text{distance}}}}{{{\text{time}}}} = \frac{{100}}{5} = ...miles/\sec \]

OpenStudy (anonymous):

20

OpenStudy (michele_laino):

correct! it is 20 miles/seconds

OpenStudy (anonymous):

????

OpenStudy (michele_laino):

we can change unit of measure, namely we can measure that speed in Km/hour

OpenStudy (anonymous):

so 20 is the right answer not what they said?

OpenStudy (michele_laino):

20 miles /seconds, is your speed, and as you can see , the corresponding unit of measure is miles/seconds

OpenStudy (michele_laino):

nevertheless, if we want to change unit of measure, of course, we get a different answer

OpenStudy (anonymous):

20

OpenStudy (anonymous):

so thats all i had to do

OpenStudy (anonymous):

@Buttercup214 its 20

OpenStudy (michele_laino):

yes! if you want the speed in miles/seconds, that is all!

OpenStudy (anonymous):

oh sorry i forgot to say thank you.